# Market Reaction to News Flows in Supply Chain Networks

### v1

@@ Methods
| Dependent Variable: | Daily percentage change of the stock price of the firm mentioned in each news article in the pre- or post-news period | ||||||
| Model: | (1) | (2) | (3) | (4) | (5) | (6) | (7) |
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| Market-specific trend | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| Sector fixed effects | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Number of observations | 9,286,677 | 9,334,585 | 9,354,107 | 9,369,523 | 9,402,470 | 9,381,512 | 9,314,442 |
@@
In column (1) of Table 3, the estimated is 0.419 and highly significant, indicating that the level of the positiveness in a news article about a firm is positively correlated with the percentage change in the firm’s stock price for 2 days before the disclosure of the news. Specifically, because the positiveness index, in equation (3), ranges from 0 to 1 (Figure 4), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.419% points higher than that of a firm mentioned most negatively () before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive and negative information included in news articles diffuses to stock markets partially through informal channels before the formal disclosure of news, generating the pre-news effect. Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms negatively mentioned, leading to endogeneity biases.
@@
Column (1) of Table 3 further finds that the estimated is 0.505, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.505% points than that of a firm mentioned most negatively after the news disclosure. The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0. In column (1), the value is close to 0, and thus we reject the null hypothesis. This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned negatively even before the disclosure of news, positive news raises stock prices of former firms more after its disclosure.
@@
These results in column (1), or for the time window of 2 days, generally hold in columns (2)-(5) for time windows of 3-5 and 30 days, while the post- and pre-news effects and the difference between the two are smaller as the time window becomes wider. By contrast, the difference between the post- and pre-news effects is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and negatively by news articles is negligible. Therefore, the positive and significant difference between the post- and pre-news effects in the short run (time windoes from 2 to 30 days) is mostly due to information diffusion through informal channels before the disclosure of news.
@@
Figure 6 graphically demonstrates the argument above. The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news effect for different time windows, whereas the right half shows the post-news effect. Although the confidence intervals are added in the figure, they are too small to be recognized. The estimated coefficients show an inverted-U shape where the coefficients are close to 0 on the left and right edges. This shape suggests the followings. The change rate of the stock price of firms positively mentioned by news articles is not intrinsically different from that of firms negatively mentioned long before the disclosure of the news. However, the change rate of the stock price of positively-mentioned firms started to increase compared with negatively-mentioned firms 30 days before the disclosure and reaches the peak 2 days after the disclosure. Then, the difference between the change rates of the stock prices of positively- and negatively-mentioned firms becomes smaller over time and converges to almost 0 180 days after the disclosure.
@@
Indirect effect on suppliers through global supply chains
@@
Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and customers through global supply chains by estimating equation 3 where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or customer of firm .
@@
Figure 7 is the graphical presentation of the results for the effect of news about firms on their suppliers, whereas Table 4 particularly shows the difference between the post- and pre-news effects and its values. The results shown in the figure and table are similar to the results for the direct effect on own firms’ stock prices (Figure 6 and Table 3). The post- and pre-news effects are positive and significant, and their difference is positive and significant when the time window is narrow. These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news through informal information diffusion and raises them more after the disclosure.
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| value | 0.00 | 0.00 | 0.23 | 0.00 | 0.00 | 0.03 | 0.00 |
@@
However, there are two notable differences between the direct effect on own firms and the indirect effect on suppliers. First, the post- and pre-news effect on suppliers and their difference is substantially smaller than the direct effects and their difference. For example, the pre-news effect on suppliers for the time window of 2 days is 0.0238 (Figure 7), while the corresponding direct effect is 0.419 (Figure 6), more than 10 times larger. Similarly, the net effect of the disclosure of news, represented by and shown in Table 4, is 0.0033 when the time window is 2 days, while the corresponding value for the direct effect is 0.086 (Table 3). Therefore, we conclude that although positive news articles about firms affect stock prices of their suppliers, the effect on suppliers is less than 10% of the direct effect on own firms. Second, the difference between the post- and pre-news effects on suppliers is positive and significant only when the time window is 2 or 3 days, although the significantly positive difference between the post- and pre-news direct effects is observed in a longer run, up to the time window of 30 days. This implies that the net effect of the disclosure of news on suppliers lasts for only several days, a very short run, after the disclosure.
Indirect effect on customers through global supply chains
Similarly, we examine the effect on customer firms and show the results in Figure 8 and Table 5. In this analysis, although the pre- and post-news effects on customers show an inverted-U shape (panel [A] of Figure 8), as in the previous results, we do not find any positive and significant difference between the post- and pre-news effects (the upper part of Table 5). Therefore, in addition to the benchmark index of positiveness of each news article using the probabilities of positive and negative sentiment from the FinBERT (Figure 4), we employ an alternative positiveness index based on only the probability of positive sentiment shown by pink bars in Figure 3. As is clear from the comparison between Figures 3 and 4, the alternative index tends to evaluate the positiveness of news articles less than the benchmark index. Thus, using the alternative index, we can examine the effect of highly positive news. Then, we find a positive and significant difference between the post- and pre-news effects, or the net effect of the disclosure of highly positive news articles, on customers for the time windows of 2 and 3 days, as shown in the lower rows of Table 5. As in the case of the effect on suppliers, the net effect on customers is small and lasts only for several days.
| 2 | 3 | 4 | 5 | 30 | 180 | 365 | |
| Benchmark positiveness index using the positive and negative probabilities of news articles | |||||||
| 0.0008 | -0.0007 | ||||||
| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Alternative positiveness index using only the positive probability of news articles | |||||||
| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
@@
We further use another sample that focuses on Japanese listed firms. The results for the direct effect of positive news articles on Japanese firms’ own stock prices shown in Figure 9 are qualitatively the same as and quantitatively similar to those using the sample of firms across the world shown in Figure 6.
@@
In addition, we examine the indirect effect of news articles about firms on stock prices of their suppliers and customers. A notable difference between this analysis on Japanese firms and the previous one on firms across the world is that more detailed domestic supply chains are identified based on firm-level surveys in the former than in the latter that mostly relies on financial statements (Data Section). Therefore, the average number of supply chain links for each Japanese firm in 2016 is 6.4, while that for each firm in the world is 3.3 (Tables 1 and 2).
The results are presented in Figure 10 and Table 6. Panels (A) and (B) of Figure 10 for the effect on suppliers and customers, respectively, show several characteristics similar to each other but different from the previous results. First, although the pre- and post-news effect on stock prices of both suppliers and customers are mostly positive and significantly different from 0, as in the previous results, the pre-news effect is negative when the time window is 180 and 365 days, unlike the previous results. The negative pre-news effect in the long run implies that stock prices of firms positively mentioned in news do not necessarily grow faster long time before the news. In other words, the effect of news is less likely to be overestimated due to selection of potentially fast-growing firms.
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| Effect on suppliers | |||||||
| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Effect on customers | |||||||
| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
Second, the pre-news effect is positive and significant, implying diffusion of information before the disclosure of news that is consistent with the previous results. However, the difference between the post- and pre-news effects that can be interpreted as the net effect of the disclosure of news is negative in the short run, or in the time window of 2-5 days and turns to be positive in the long run (Table 6). This result is completely opposite to the previous ones using data for firms across countries (Tables 4 and 5). In other words, positive information about a firm diffuses through informal channels and raises the change rate of stock prices of the firm’s suppliers and customers before the information is disclosed in a news article. However, the disclosure of the information by a news article lowers the change rate of stock prices compared with that in the pre-news period, although the post-news effect is still positive. One possible reason for this is that investors who can receive information about firms through informal channels are more knowledgeable and recognize diffusion of performance through supply chains more accurately than other investors who rely on formal news. Therefore, once the knowledgeable investors receive positive information about a firm before the disclosure by news, they invest in the firm’s suppliers and customers. However, other investors react less to positive news about firms’ suppliers and customers.
Finally, the pre- and post-news effects on suppliers and clients found in the sample of Japanese firms are larger than those in the sample of firms across the world. This finding, combined with the previous interpretation about knowledgeable investors, suggests that supply chain ties between suppliers and customers in Japan, known as keiretsu, are often stronger than in other countries, because they are often associated with shareholding relationships, information exchanges, technical assistance and research collaboration.[43, 44] Therefore, diffusion of performance through supply chains may be more prominent in Japan than in others, and this possibility is recognized by at least knowledgeable investors in Japan.
Conclusion
@@
This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys.
@@
We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels. Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion. However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
@@ Acknowledgements
The authors thank Yuito Nishi for his technical contribution.

### v3

@@ 4 Results and Discussion
| Dependent Variable: | Daily percentage change of the stock price of the firm mentioned in each news article in the pre- or post-news period | |||||||
| Model: | (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
| Time window (days) [] | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| Market-specific trend | Yes | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| Sector fixed effects | Yes | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Number of observations | 9,409,978 | 9,543,586 | 9,562,439 | 9,565,054 | 9,562,897 | 9,460,176 | 8,986,209 | 8,546,395 |
@@
In column (1) of Table 4, the estimated is 0.323 and highly significant, indicating that the probability of positive sentiment in a news article about a firm is positively correlated with the percentage change in the firm’s stock price from 2 days to 1 day before the disclosure of the news. Since the positiveness measure, in equation (3), ranges from 0 to 1 (Figure 3), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.323% points higher than that of a firm mentioned with no positive sentiment () before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive information included in news articles diffuses to stock markets partially through private channels before the formal disclosure of news, generating the pre-news effect. Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms with no positive mention, leading to endogeneity biases.
@@
Column (1) of Table 4 further finds that the estimated is 0.923, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.923% points than that of a firm mentioned with no positive sentiment after the news disclosure. The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0. In column (1), the value is close to 0, and thus we reject the null hypothesis. This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned with no positive sentiment even before the disclosure of news, positive news raises stock prices of former firms more by 0.6% points after its disclosure.
@@
These results in column (1) for the time window of 1 day generally hold in columns (2)-(5) for time windows of 2-5 and 30 days, while the post- and pre-news effects and the difference between the two decrease as the time window becomes wider. By contrast, the difference between the post- and pre-news effects is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and those with no positive mention in news articles is negligible.
@@
Panel (A) of Figure 5 graphically demonstrates the argument above. The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news effect for different time windows, whereas the right half shows the post-news effect. Although the confidence intervals are added in the figure, they are too small to be recognized. The estimated coefficients show an inverted-U shape where the coefficients are close to 0 on the left and right edges. This shape suggests the followings. The change rate of the stock price of firms positively mentioned by news articles is not intrinsically different from that of firms negatively mentioned long before the disclosure of the news. However, the change rate of the stock price of positively mentioned firms began to rise compared to firms with no positive mention 30 days before the disclosure, peaking one day after the disclosure. Then, the difference in the change rates of stock prices between positively mentioned firms and those with no positive mention gradually decreases over time, converging to nearly zero 180 days after the disclosure.
The effect of the probability of negative sentiment in each news article shown in Panel (B) of Figure 5 is consistent with the effect of the probability of positive sentiment in Panel (A). The effect of negative sentiment in a news article is negligible 180 days or more before its disclosure but becomes negative and significant 30 days or fewer before the disclosure, implying information diffusion through private channels. After the disclosure, the negative effect becomes larger in the absolute term, suggesting a negative net effect of the disclosure of negative news on stock prices, in addition to the negative effect through private channels.
@@
Indirect effect on suppliers and clients through global supply chains
@@
Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and clients through global supply chains by estimating equation (3) where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or client of firm .
@@
Panels (A) and (B) of Figure 6 are the graphical presentation of the results for the effect of positive and negative news about firms, respectively, on their suppliers. Table 5 particularly shows the difference between the post- and pre-news effects of positive news and its values. The corresponding table for the effect of negative news is omitted here but is available in the Supplementary Information (SI Table 1). The results shown in the figure and table are similar to the results for the direct effect on own firms’ stock prices (Figure 5 and Table 4). The post- and pre-news effects of positive news are positive and significant, and their difference is positive and significant when the time window is narrow. These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news due to information diffusion through private channels and raises them more after the disclosure due to the additional effect of news medias. The effect of negative news articles is opposite in the direction and similar in size to the effect of positive ones, indicating the same conclusion for negative news.
| Time window (days) | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.27 | 0.00 | 0.00 |
@@
However, there are two notable differences between the direct effect on own firms and the indirect effect on suppliers. First, the post- and pre-news effect on suppliers and their difference is substantially smaller than the direct effects and their difference. For example, the pre-news effect on suppliers for the time window of 1 days is 0.0293 (Figure 6), while the corresponding direct effect is 0.323 (Figure 5), more than 10 times larger. Similarly, the net effect of the disclosure of news, represented by and shown in Table 5, is 0.033 when the time window is 1 days, while the corresponding value for the direct effect is 0.600 (Table 4). Therefore, we conclude that although positive news articles about firms affect stock prices of their suppliers, the effect on suppliers is less than 10% of the direct effect on own firms. Second, the difference between the post- and pre-news effects on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, meaning that the effect of the disclosure of positive news about a firm on its suppliers shrinks by 88% in 5 days. By contrast, its effect on the firm’s own stock price shrinks from 0.600 to 0.157 by 74%. This comparison implies that the net effect of the disclosure of news on suppliers lasts shorter after the disclosure than the effect on own stock prices.
Similarly, we examine the effect of positive and negative news articles on stock prices of client firms and illustrate the results in Panels (A) and (B) of Figure 7, respectively. Table 6 additionally presents the net effect of positive news, while the corresponding table for negative news can be found in the Supplementary Information (SI Table 2). These results are quite similar to those of the effect on suppliers shown in Figure 6 and Table 5. In summary, we find a positive (negative) and significant pre- and post-news effect of positive (negative) news and a positive (negative) and significant net effect of the disclosure of news. However, the effect of news about a firm on its clients is substantially smaller and lasts shorter than the effect on its own.
| Time window (days) | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.13 | 0.00 | 0.00 |
@@
We further use another sample that focuses on Japanese listed firms. The results for the direct effect of positive and negative news articles on Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5. One subtle difference is that in Figure 8 for Japan, the post-news effect of negative news using the time window of 2 days is larger in the absolute term than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite. This finding suggests that the market reaction to negative news in Japan is slower than that to positive news in Japan and to positive and negative news in the world.
@@
In addition, we examine the indirect effect of news articles about Japanese firms on stock prices of their suppliers and clients. A notable difference between this analysis on Japanese firms and the previous one on firms across the world is that more detailed domestic supply chains are identified based on firm-level surveys in the former than in the latter that mostly relies on financial statements (Data Section). Therefore, the average number of supply chain links for each Japanese firm in 2016 is 6.4, while that for each firm in the world is 3.3 (Tables 1 and 2).
The results for the effect of positive and negative news on suppliers are demonstrated in Panels (A) and (B) of Figure 9, respectively, showing three notable differences from the previous results in Figure 6. First, the pre-news effect due to information diffusion through private channels is generally larger for Japanese firms than for others. For example, using the time window of 1 and 2 days, the pre-news effect of positive news on suppliers is 0.0090 and 0.0079 for Japan but 0.0029 and 0.0017 for others, respectively. Similarly, the post-news effect of negative news using the same windows is -0.0089 and -0.0066 for Japan and -0.0046 and -0.0025 for others, respectively.
Second, the post-news effect of positive news is smaller than its pre-news effect (Panel [A] of Figure 9), although the post-news effect is always larger in size than the pre-news effect in the previous results. This finding implies that although positive information about a firm tends to diffuse actively through private channels, such as suggestions of financial companies, SNS, and word of mouth, before its disclosure by a news article and increase stock prices of suppliers of the firm. However, once the positive information is disclosed in a news article, the growth rate of the suppliers’ stock prices declines compared with that in the pre-news period. In other words, the net effect of the disclosure of positive news about a Japanese on stock prices of its suppliers is negative, although the previous results found the opposite.
Finally, we do observe that the post-news effect is larger than the pre-news effect in the case of negative news, as in the previous results (Panel [B] of Figure 9). However, the pre-news effect of negative news for the time window of 1 day is smaller in the absolute term than that for the longer time window of 2-5 days, although the previous result for firms across the world showed the opposite (Panel [B] of Figure 6). This, combined with the large negative post-news effect for the time window of 1 day, implies that the disclosure of negative information about a firm by a news article affects stock prices of its suppliers more substantially than diffusion of the information through private channels.
We further estimate the effect on clients and show the results in Figure 10. Compared with the previous results for firms across the world shown in Figure 7, we find that the three notable differences found in the effect on suppliers still hold in the effect on clients: (1) The average effect is generally larger for Japanese firms than for firms across the world; (2) The net effect of the disclosure of positive news about a firm on stock prices of its clients is negative; (3) The pre-news effect of negative news for the time window of 1 day is smaller than that for longer time windows.
In summary, we find that the way positive and negative information about firms affects stock prices of their suppliers and clients before and after the disclosure of the information by news media is different between firms in Japan and those in other countries. Specifically, we find that positive news propagates more before its disclosure than after in the Japanese data. Although we cannot provide evidence for any mechanism behind the differences, one possible reason is the difference in the quality of supply chain information. The supply chain data used for firms worldwide in this study relies on financial statements, whereas the data for Japan is based on a questionnaire-based survey. Although the former is automatically determined by criteria based on sales volume, the latter identifies important trade partners, which are not determined only by sales volume. Since investors who can receive information about firms through private channels are more knowledgeable about supply chains and recognize the diffusion of performance through supply chains more accurately than other investors who rely on news media, once the knowledgeable investors receive positive information about a firm before its disclosure by news media, they invest in the firm’s suppliers and clients. This results in a large increase in stock prices of the suppliers and clients before the disclosure of the news. However, other investors who rely on news media and do not understand the role of supply chains react less strongly to positive news after its disclosure.
Another possible reason for this is the distinct nature of Japan’s supply chains, known as keiretsu. Links between suppliers and clients in these relationships are often stronger than those in other countries, as they are frequently associated with shareholding relationships, information exchanges, technical assistance, and research collaboration [67, 68]. These strong relationships may explain why positive news about firms may have a greater impact on their suppliers and clients. Because investors recognize the strength of these ties and anticipate the transmission of performance through supply chains, news about firms tends to have a greater impact on the stock prices of their suppliers and clients in Japan compared to other countries. However, these two reasons cannot explain why negative news does not propagate before disclosure, which should be investigated in future research.
5 Conclusion
@@
This study examines how positive and negative news about firms affects the change rate of their stock prices and, moreover, how it affects the change rate of stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements and supplemented by other information on the website and news medias, while those of Japanese firms are taken from large-scale firm-level surveys.
@@
We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through private channels, such as SNS and word of mouth. Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels. However, the post-news effect on suppliers and clients in Japan can be smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
The important opening question not addressed by this study is the significant heterogeneity of supply chain relationships, which can lead to varying impacts of news sentiment. For example, some suppliers are easily replaceable, while others are not; some trade relationships are long-term, potentially reflecting trust or sunk costs; additional links, such as shareholding or dual board membership, may exist; and the recognition of supply chain relationships by investors can vary. Examining the heterogeneity presents an important direction for future research.
Data availability
The data that support the findings of this study are available from FactSet Research Systems Inc. and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics. but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with permission of FactSet Research Systems Inc. and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics.
@@ Acknowledgements
The authors thank Wataru Souma and Takayuki Yumoto for valuable discussions and Yuito Nishi for his technical contribution.